OpenStudy (anonymous):

Medal for help :) Suppose y varies directly with x and y =25 when x =140.What is the value of x when y =36? a.140 b.165 c.201.6 d.176

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

@am!rah

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (ibbutibbu.):

this is a problem of propotions

OpenStudy (phi):

Suppose y varies directly with x when you see that you should think: y = k x where k is a constant you need to find

OpenStudy (anonymous):

Ok

OpenStudy (phi):

to find k, you use the info they gave you y =25 when x =140 use those numbers in y = k x what do you get ?

OpenStudy (anonymous):

Oh B

OpenStudy (phi):

it's a bit silly to guess. You have to go through the steps, and the answers comes out at the end y =25 when x =140 use those numbers in y = k x in other words, you get 25= k*140 divide both sides by 140 to get 25/140 = k now you can answer your question

OpenStudy (phi):

What is the value of x when y =36 y = (25/140) x replace y with 36 36 = (25/140)x multiply both sides by 140/25 to get 140*36/25 = x

OpenStudy (phi):

is your calculator broken? what is 140*36/25?

OpenStudy (anonymous):

Oh Ish 201.6 ty that one got me confused